OpenStudy (anonymous):

what is the y-intercept of this line? y+4=-4(x-2) can anyone help me!!

OpenStudy (mertsj):

Make the equation look like this: y = mx + b b will be the y intercept

OpenStudy (anonymous):

\[y+4 = -4(x-2)\]\[y +4 = -4x + 8\]\[y= -4x +4\] Y = Mx+b

OpenStudy (anonymous):

Another way to understand this is that the Y-Intercept of a line is that point where it cuts the Y-Axis. This means that the value of X is Zero at the Y-Intercept. So all you do is substitute x=0 in the equation and solve for Y. In this case the it is quite easy to see that Y=4 when X=0, and thus the Y-intercept is (0,4). Also note that this is true for ANY curve - not necessarily only a line.